I need help with the addition method. I have a problem that i cannot figure out. Please help.
6x-5y+z=2
3x-4y+z=2
4x+2y-z=-5

The available ansswers are as follows:
A. (-3, -9, -25)
B. (0, -1, -3)
C. real numbers
D. No soulution
Thank you

Answer by venugopalramana(3286)

I need help with the addition method. I have a problem that i cannot figure out. Please help.
6x-5y+z=2.............................I
3x-4y+z=2............................II
4x+2y-z=-5...............................III
EQN.I-EQN.II
3X-Y=0........................................IV
EQN.II+EQN.III
7X-2Y=-3.....................................V
2*EQN.IV-EQN.V
-X=3...X=-3
Y=3X=-9
6*-3-5*(-9)+Z=2
Z=-25

A IS THE BANSWER

The available ansswers are as follows:
A. (-3, -9, -25)
B. (0, -1, -3)
C. real numbers
D. No soulution